OpenStudy (tiffanymak1996):

Let f(x,y) =(x^3 +y^3) ^(1/3) At what points if any, does fy(x,y) fail to exist?

OpenStudy (anonymous):

\(f_y(x,y)=y^2(x^3+y^3)^{-2/3}\) so it will not exist where \(x^3+y^3=0\), or to write it other way \(y^3=-x^3\) solve this equation to get the points where this happens

OpenStudy (anonymous):

good work @myko

OpenStudy (tiffanymak1996):

in the answer of my book, equation \[y \neq0\] is also an answer along with y=-x, how do i get y does not equal 0?

OpenStudy (anonymous):

Other way to write it: y^3+x^3=(x+y)(x^2-xy+y^2)=0 so you see that x+y=0 makes the hole thing to be 0, right? so y=-x is a place where \(f_y\) will not exist, and it's a line trough the origin

OpenStudy (anonymous):

ty @oldrin.bataku

OpenStudy (tiffanymak1996):

well, i can also work that out from y^3 = -x^3 , but what about y does not equal 0?

OpenStudy (anonymous):

\(y \neq 0\) by it self is not a place where \(f_y\) not exist.

OpenStudy (anonymous):

maybe what your book means is that if \(y \neq 0\) then it have to be on the y=-x line

OpenStudy (tiffanymak1996):

oh, ok, thx!